A 13 foot ladder is positioned against a wall so that the base of the ladder is 5 feet from the wall. How high is the top of the ladder above the ground?
Answer by TimothyLamb(4379)

h = sqrt( aa + bb )
13 = sqrt( aa + 5*5 )
13^2 = ( sqrt( aa + 5*5 ) )^2
169 = aa + 25
aa = 169 - 25
a = sqrt( 169 - 25 )
a = 12
---
answer:
12 feet above the ground
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---